OpenStudy (anonymous):

A machinist turns the power on to a grinding wheel, which is at rest at time t = 0.00 s. The wheel accelerates uniformly for 10 s and reaches the operating angular velocity of 25 rad/s. The wheel is run at that angular velocity for 37 s and then power is shut off. The wheel decelerates uniformly at 1.5 rad/s2 until the wheel stops. In this situation, the time interval of angular deceleration (slowing down) is closest to: A) 17 s B) 15 s C) 19 s D) 21 s E) 23 s

OpenStudy (anonymous):

it really close to 23s in not really sure but i did calculation in my head and i trust it right lol

OpenStudy (anonymous):

can you provide your steps, thanks

OpenStudy (anonymous):

\[\alpha = \frac{ d \omega }{ dt }\]Angular acceleration is derivative of angular velocity the same way as linear acceleration is derivative of linear velocity \[a = \frac{ dv }{ dt}\]

OpenStudy (anonymous):

To make I'm on the same page with you I took these steps t = (25 rad/s) / 1.5 rad/s^2 = 16.6 ~ 17s